We wish to estimate what percent of adult residents in a certain county are parents. Out of 200 adult residents sampled, 90 had kids. Based on this, construct a 95% confidence interval for the proportion p of adult residents who are parents in this county.

Express your answer in tri-inequality form. Give your answers as decimals, to three places.

< p < Express the same answer using the point estimate and margin of error. Give your answers as decimals, to three places.

p = ±±
Answer #1

Solution :

Given that,

n = 200

x = 90

Point estimate = sample proportion = \hat p = x / n = 90 / 200 = 0.45

1 - \hat p = 1 - 0.45 = 0.55

At 95% confidence level

\alpha = 1 - 95%

\alpha =1 - 0.95 =0.05

\alpha/2 = 0.025

Z\alpha/2 = Z0.025 = 1.960

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.96 (\sqrt((0.45 * 0.55) / 200)

= 0.069

A 95% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.45 - 0.069 < p < 0.45 + 0.069

( 0.381 < p < 0.519 )